# A Temporal Convolutional Network-Based Approach and a Benchmark Dataset for Colonoscopy Video Temporal Segmentation

### v1

@@ 2.1 Annotation of the REAL-Colon dataset
For this study, all 60 videos in the REAL-Colon dataset underwent additional annotation. Eight medical image annotation specialists, supervised by an expert gastroenterologist from a pool of three experts, followed a standardized process using a specialized in-house annotation tool capable of timestamp placement. The process included multiple feedback rounds and can be summarized as follows. First, three colonoscopy phase labels — outside, insertion, and withdrawal — were annotated, each with a start and stop frames. Instruction was to make the withdrawal phase start when the endoscope reached the appendiceal orifice or ileocecal valve, and the endoscopist initiated cecum exploration following cleaning. The transition between the colon interior and exterior was marked at the frame where the anus was entered or exited. Then, the withdrawal phase was annotated into seven colon segment sub-classes: ceacum, ileum, ascending colon, transverse colon, descending colon, sigmoid colon, and rectum. The ascending, transverse, and descending colon were delineated by identifying the first frame in which the hepatic and splenic flexure was visible. Similarly, the cecum was defined by identifying the ileocecal valve and appendiceal orifice, while the rectum was defined by identifying the rectosigmoid junction. In cases where the endoscope entered the ileocecal valve and explored the terminal ileum, distinguished by its specific mucosal appearance, the corresponding label was assigned to those frames.
@@ 4 Discussion
A key contribution of this research was the effective adaptation of TCN-based models to address this problem. In particular, we propose ColonTCN, a specialized TCN architecture characterized by multiple custom TBs forming a deep hierarchy. These TBs, through the exploitation of double dilated residual convolutions and regularization, enable frame-by-frame contextual understanding while efficiently achieving a sufficient temporal receptive field to comprehend the entire procedure. Two categories of models were compared against ColonTCN: ASFormer and TCNs and MS-TCNs from TeCNO [32, 5]. The latter were chosen as the most advanced TCN-based architectures in surgical workflow analysis, which do not rely on domain-specific modifications. The former was selected because it is a modern, lightweight transformer-based model and was used in similar research [21]. From the results reported in Table 8 it can be seen how both class of methods struggle to achieve an adequate receptive field without significantly increasing their their total number of parameters. In the case of ASFormer, despite using windowed attention, it is computationally expensive to model such long videos due to the inherently high computational cost of attention mechanisms, which scale quadratically with the length of the input sequence. In the case of models from TeCNO, these models require more temporal convolution layers than ColonTCN to achieve a comprehensive receptive field, which can hinder model learning by adversely affecting gradient flow. The TBs proposed for ColonTCN have shown to better and more efficiently learn from colonoscopy video data.

### v2

@@ 2.1 Annotation of the REAL-Colon dataset
For this study, all 60 videos in the REAL-Colon dataset underwent additional annotation. Eight medical image annotation specialists, supervised by an expert gastroenterologist from a pool of three experts, followed a standardized process using a specialized in-house annotation tool capable of timestamp placement. The process included multiple feedback rounds and can be summarized as follows. First, three colonoscopy phase labels — outside, insertion, and withdrawal — were annotated, each with a start and stop frame. Instruction was to make the withdrawal phase start when the endoscope reached the appendiceal orifice or ileocecal valve, and the endoscopist initiated cecum exploration following cleaning. The transition between the colon interior and exterior was marked at the frame where the anus was entered or exited. Then, the withdrawal phase was annotated into seven colon segment sub-classes: ceacum, ileum, ascending colon, transverse colon, descending colon, sigmoid colon, and rectum. The ascending, transverse, and descending colon were delineated by identifying the first frame in which the hepatic and splenic flexure was visible. Similarly, the cecum was defined by identifying the ileocecal valve and appendiceal orifice, while the rectum was defined by identifying the rectosigmoid junction. In cases where the endoscope entered the ileocecal valve and explored the terminal ileum, distinguished by its specific mucosal appearance, the corresponding label was assigned to those frames.
@@ 4 Discussion
A key contribution of this research was the effective adaptation of TCN-based models to address this problem. In particular, we propose ColonTCN, a specialized TCN architecture characterized by multiple custom TBs forming a deep hierarchy. These TBs, through the exploitation of double dilated residual convolutions and regularization, enable frame-by-frame contextual understanding while efficiently achieving a sufficient temporal receptive field to comprehend the entire procedure. Two categories of models were compared against ColonTCN: ASFormer and TCNs and MS-TCNs from TeCNO [32, 5]. The latter were chosen as the most advanced TCN-based architectures in surgical workflow analysis, which do not rely on domain-specific modifications. The former was selected because it is a modern, lightweight transformer-based model and was used in similar research [21]. From the results reported in Table 8 it can be seen how both class of methods struggle to achieve an adequate receptive field without significantly increasing their total number of parameters. In the case of ASFormer, despite using windowed attention, it is computationally expensive to model such long videos due to the inherently high computational cost of attention mechanisms, which scale quadratically with the length of the input sequence. In the case of models from TeCNO, these models require more temporal convolution layers than ColonTCN to achieve a comprehensive receptive field, which can hinder model learning by adversely affecting gradient flow. The TBs proposed for ColonTCN have shown to better and more efficiently learn from colonoscopy video data.